Gucchi:

Which of the following statements best describes the effect of replacing the graph of y = f(x) with the graph of y = f(x) − 11? The graph of y = f(x) will shift up 11 units. The graph of y = f(x) will shift right 11 units. The graph of y = f(x) will shift left 11 units. The graph of y = f(x) will shift down 11 units.

Gucchi:

Am i right with D?

Gucchi:

@axie

Gucchi:

@tetsxpreme

Gucchi:

@supie

CheeseStick7793:

the answer is The graph of y = f(x) will shift down 11 units.

CheeseStick7793:

happy to help yes u were correct

CheeseStick7793:

:)

Gucchi:

@cheesestick7793 wrote:
the answer is The graph of y = f(x) will shift down 11 units.
explanation pls?

Gucchi:

how am i right

CheeseStick7793:

The parent function is y = f(x) The transformation rule is: If the parent function f(x) is translated 'a' units down then the equation of the parent function become f(x)-a. Here the equation of the translated function is y = f(x) - 11. It means that a = 11 And from the above rule, the function f(x) is translated 11 units down. D is correct.

Gucchi:

tet do you agree?

TETSXPREME:

Yes C:

TETSXPREME:

Good Job

CheeseStick7793:

mhm

Gucchi:

ty